### v1

# Federated Foundation Models in Harsh Wireless Environments: Prospects, Challenges, and Future Directions

Abstract
Foundation models (FMs) have shown remarkable capabilities in generalized intelligence, multimodal understanding, and adaptive learning across a wide range of domains. However, their deployment in harsh or austere environments – characterized by intermittent connectivity, limited computation, noisy data, and dynamically changing network topologies – remains an open challenge. Existing distributed learning methods such as federated learning (FL) struggle to adapt in such settings due to their reliance on stable infrastructure, synchronized updates, and resource-intensive training. In this work, we explore the potential of Federated Foundation Models (FFMs) as a promising paradigm to address these limitations. By integrating the scalability and generalization power of FMs with novel decentralized, communication-aware FL frameworks, we aim to enable robust, energy-efficient, and adaptive intelligence in extreme and adversarial conditions. We present a detailed breakdown of system-level constraints in harsh environments, and discuss the open research challenges in communication design, model robustness, and energy-efficient personalization for these unique settings.
Index Terms:
Foundation Models, Federated Learning, Harsh Environments
I Introduction
Machine Learning (ML) has revolutionized numerous domains, from healthcare and smart cities to transportation and environmental monitoring [1]. In their conventional forms, ML models are trained and deployed to handle a single modality of data (e.g., image, text) and a single downstream task (e.g., classification of a particular dataset). However, intelligent services offered in modern wireless networks are increasingly demanding the simultaneous processing of multiple modalities of data collected at wireless devices, and handling the execution of a variety of downstream tasks. For example, in smart driving cars, image, LiDAR, and sensor readings are jointly leveraged to support tasks such as collision avoidance, traffic sign recognition, and driver assistance.
We have witnessed a similar trend towards multi-task and multi-modality in the ML area. The community is migrating towards ML models capable of conducting multiple ML tasks through the use of various input data modalities. The advent of Large Language Models (LLMs) was a giant leap in this direction, as LLMs are capable of executing multiple language/text-related tasks (e.g., code synthesis, text generation, summarization, etc.). More recently, the need for ML models with larger scopes led to another major innovation in the ML area: Foundation models (FMs). FMs enable the use of simultaneous modalities of data (through dedicated encoders), such as text, audio, image, and video, to execute multiple downstream tasks (through dedicated task heads), such as audio generation, translation, image generation, video generation, and text refinement. Further, through their unified architecture, FMs can enable knowledge sharing across multiple modalities, transfer knowledge across tasks, and swiftly adapt to unseen data through lightweight Parameter Efficient Fine-Tuning (PEFT). Popular PEFT techniques include adapters, Low-Rank Adaptations (LoRA), and prompt tuning [2]. Notably, their strong zero/few-shot learning capabilities have made PEFT particularly appealing for on-device fine-tuning and inference, especially when considering resource-constrained devices.
Figure 1: Multiple levels of deployment challenges arise when aiming to conduct Federated Learning (FL) in harsh environments, including unreliable communication layers and noisy, incomplete local data for training which create bottlenecks for server-side coordination.
Yet, due to their nascent nature, FMs have not yet been deployed across networks with unusual or extreme requirements, which the communication systems community collectively refers to as harsh or austere wireless networks/environments. In particular, in this work, for the first time in the literature, we shed light on the deployment of FMs in harsh environments. Harsh environments (encountered in e.g., undersea exploration, underground mines, polar regions) often yield data that are sparse, corrupted by sensor failures/miscalibration, or biased toward narrow operating regimes. ML models trained solely on these local datasets will thus tend to converge to suboptimal solutions and perform poorly when exposed to changing scenarios. This calls for unlocking the access of FMs to distributed data collected across multiple harsh environment regions during their fine-tuning/training stage. Given well-known challenges in centralizing large volumes of data collected across these regions, doing so requires fine-tuning/training according to distributed learning techniques.
Federated Learning (FL) [3] is a popular technique for distributed ML which allows multiple devices to collaboratively train a shared global model without exchanging raw data, thereby preserving privacy and reducing communication overhead. Consequently, it is compelling to consider leveraging FL-driven fine-tuning/training of FMs, referred to as Federated Foundation Models (FFMs) [4], in harsh environments. However, there are only a few existing works on FFMs [5], and the literature is yet to explore them in harsh environments, motivating this study. In particular, the operational realities of such environments, ranging from intermittent connectivity and physical hazards to extreme resource constraints and rapidly shifting data distributions (depicted in Fig. 1), demand a careful rethinking of how FFMs are implemented in these environments [6, 7].
Summary of contributions. In this work, our contributions are as follows:
  • We highlight the advantages of FMs for providing intelligence in harsh environments, outlining their capabilities for multimodal perception, rapid adaptability, task versatility, and robustness to noise and distribution shifts.
  • We identify the unique challenges that arise when deploying FFMs in harsh environments, including the impacts of fragile and failure-prone devices, asynchronous and stale model updates, and the difficulty of maintaining or replacing nodes in inaccessible locations.
  • We outline open research directions for realizing FFMs in harsh environments, including multimodal alignment, drift mitigation, robustness to unreliable communication and corrupted data, and energy-efficient operation.
Figure 2: Illustration of a foundation model architecture. Multimodal inputs are processed by a shared LLM decoder. Parameter-efficient tuning methods, including adapters and low-rank adaptation, enable lightweight specialization. The unified representation supports diverse downstream tasks.
Figure 3: Foundation Models (FMs) have the potential to enable robust, multimodal, and adaptable intelligence across heterogeneous harsh environments, with resilience to noise and distribution shifts. Combined with Federated Learning (FL), FFMs empower edge devices to collaboratively perceive, reason, and act under extreme physical and communication constraints.
II Key Concepts of FFMs and Harsh Environments
II-A Foundation Models: Structure and Key Advantages
FMs are large-scale machine learning architectures trained on massive and diverse datasets to capture general-purpose representations across modalities and tasks. We illustrate an FM architecture in Fig. 2. FMs are typically built on deep encoder-decoder or decoder-only backbones, which learn unified embedding spaces capable of representing multimodal signals. Unlike conventional models trained from scratch for each application, FMs are designed to be adapted rather than fully retrained by using aforementioned PEFT techniques. Moreover, FMs can function as modular agents, where encoders extract shared semantic representations from diverse inputs, and downstream task heads or lightweight adapters customize the model to specific objectives.
Several unique advantages of FMs naturally emerge from these properties, which we outline in Fig. 3. First, their multimodal perception allows seamless fusion of heterogeneous sensor inputs. For example, in an undersea environment, unmanned underwater vehicles (UUVs) may collect sonar, acoustic, optical, or infrared data [8]. FMs can provide a unified representation space to handle these varied modalities without retraining for every sensor configuration. Second, FMs are inherently versatile across tasks: the same backbone can support anomaly detection, classification, tracking, or signal interpretation, which is particularly valuable when a system must execute multiple functions under strict energy budgets, e.g., mining robots, underwater nodes, or forest-deployed drones. Third, FMs enable rapid adaptation through local context, e.g., a forest monitoring station needing to adapt to seasonal acoustic patterns or a satellite adapting to orbit-specific variations. With lightweight tuning, models can be personalized quickly to unique conditions. Finally, FMs exhibit robustness to distribution shifts and sensor noise. Their large-scale pretraining and reasoning abilities help maintain stable performance under heterogeneous or degraded inputs, such as noisy underwater acoustics, seasonal forest changes, or intermittent interferences in dynamic environments.
In practice, these advantages are realized through design choices that balance capability and resource use. PEFT techniques keep the number of trainable parameters small, reducing memory footprint and on-device computing cost. Sparse or mixture-of-experts backbones activate only a subset of parameters per token, lowering inference cost while preserving expressivity. Retrieval-augmented components can inject local context (e.g., mission logs, recent observations) at inference time without altering core weights, which is especially useful when connectivity is intermittent. Together, the modular architecture and parameter-efficient adaptation position FMs as a strong substrate for building systems that must perceive, reason, and act across diverse sensing conditions with limited resources.
II-B Training/Fine-Tuning of FFMs
The training/fine-tuning of FFMs are governed by iterative FL protocols, where devices repeatedly perform local updates and participate in aggregation rounds. At the beginning of training, a FM is distributed to all participating devices as the initialization. Each device then updates its local copy of the model using its private dataset, which reflects the unique sensing conditions and modalities of that environment. A central design choice is the on-device trainable subset of the model. Finetuning techniques such as adapter-only updates (e.g., bottleneck adapters or LoRA blocks inserted in attention/FFN layers) strike a favorable balance between accuracy and on-device cost; head-only tuning is lighter but can underfit multimodal shifts; partial-backbone unfreezing improves adaptability but raises memory and bandwidth requirements. For multimodal FFMs, encoders can be tuned asymmetrically (e.g., update the acoustic encoder while freezing the vision stack) and merged with a shared decoder; per-task heads are typically specialized locally and aggregated with task-aware weighting.
Local optimization often uses lightweight schedules to respect device constraints: mixed-precision training, gradient accumulation to emulate larger batches, and short-horizon fine-tuning with cosine decay after a brief warmup. Regularizers that stabilize the shared embedding, such as consistency losses between successive local checkpoints or distillation from the received global model, help curb representation drift when devices see narrow or noisy data slices. When memory is tight, activating a subset of adapters per step reduces footprint without halting progress.
Following local training, devices share their parameter changes or updated models with others according to the chosen aggregation strategy. In centralized setups, a single server collects updates, performs weighted averaging (e.g., FedAvg), and broadcasts the aggregated model back to the devices. In decentralized or serverless setup, devices exchange updates directly with peers over a peer-to-peer topology, reaching consensus without reliance on a single coordinator. In hierarchical configurations, aggregation occurs in multiple tiers, where devices transmit models through intermediate aggregators.
This cycle of local update and aggregation is repeated for multiple communication rounds as shown as Figure 4, allowing the FFM to progressively integrate knowledge from diverse environments while maintaining adaptability to local conditions. Once convergence is reached, the aggregated model can be deployed across all devices, with the same protocol continuing in an online manner to support continuous adaptation in response to new tasks, shifting data distributions, or changing network conditions.
II-C Definition of Harsh Environments
Harsh environments refer to physical and network settings where fundamental assumptions of reliable communication, abundant energy supply, and stable infrastructure no longer hold. Unlike conventional distributed learning scenarios, these environments impose compounded physical constraints (extreme pressure, temperature, salinity, humidity, vibration, or structural instability) that directly impact sensing, computation, and communication. At the same time, network-level characteristics in such environments are dominated by sparsity, intermittency, and fragility of links, often with high latency and asymmetric bandwidth. In these contexts, devices must operate with heightened autonomy, as frequent synchronization with a centralized server is infeasible. Instead, learning must be resilient to missing updates, highly resource-efficient, and capable of fusing heterogeneous multi-modal data for real-time local inference. These commonalities define the operational essence of harsh environments: they are settings where the physical world itself constrains not only communication channels but also the very feasibility of collecting, sharing, and aligning data across distributed agents.
The physical layer of communication is especially strained in such settings. Harsh environments often require reliance on lower radio frequencies, which provide stronger propagation and penetration properties than higher-frequency bands. Vintage very high frequency (VHF) ranges, though long overlooked, may be revisited in this context, supported by advances in noncoherent signaling, coding, AI-based adaptation, and interference management. In the harshest cases, where electromagnetic propagation becomes unreliable or virtually unusable, alternative modalities such as pressure-wave communication must be considered. Acoustic signaling underwater and mechanical pressure-wave propagation through pipes or boreholes are widely used in industry yet remain underexplored academically. For FFMs to function effectively, fundamental limits and channel models of these unconventional communication modes must be established, enabling protocol designs tailored to extreme conditions.
The sparsity and instability of communication in harsh environments necessitate collaborative sensing and decision-making across multiple stations or devices. In many cases, local clients must perform rapid, context-aware actions in response to evolving conditions without relying on frequent connectivity to centralized infrastructure. This elevates the need for models capable of on-site inference with lightweight adaptation.
To contextualize these requirements, we highlight four representative settings that impose distinct physical and link-layer characteristics. Underground mines exhibit severe non-line-of-sight propagation, metallic and geological interference, and the constantly evolving tunnel geometries that alter coverage needs as excavation proceeds, all under the shadow of hazards like collapses or gas leaks that can abruptly sever communication. Underwater acoustic networks face an entirely different set of difficulties: bandwidth is extremely limited, propagation delays are long and variable, Doppler spread causes signal distortion, and stratification-driven refraction bends signals in unpredictable ways. Forest deployments are hindered by dense vegetation and unpredictable weather that block and scatter signals, forcing reliance on opportunistic encounters (when drones or wildlife-tagged sensors happen to come into range). Satellites, in contrast, must contend with orbital dynamics: fast handovers, intermittent visibility windows, and stringent energy and payload limits, all of which complicate coordination across layers that may include high-altitude platforms or UAVs. These examples characterize the operational backdrop for learning systems in harsh environments. Section III-B translates these physical and link-layer realities into concrete challenges for training and inference, while Figure 5 summarizes the mapping from physical constraints to network effects.
Figure 4: Overview of FFM learning architectures. Edge devices perform local training and communicate updates through centralized, decentralized (peer-to-peer), or hierarchical topologies. Model collection leverages strategies such as weighted averaging and mixture of experts for training optimization.
III FFMs in Harsh Environments: Motivation and Core Challenges
III-A Why FFMs are needed in Harsh Environments
Foundation models bring unique capabilities, multimodal reasoning, task versatility, and rapid adaptation, that are essential for devices deployed in extreme conditions. Yet in isolation, even the most powerful FM quickly loses reliability in harsh environments. A single mining robot, undersea node, or remote satellite can only perceive a narrow and noisy slice of its environment. Left on its own, the model drifts toward biased local optima, overfits to corrupted signals, or fails to develop representations for modalities it never encounters. In practice, this means a locally adapted FM may perform well on one platform but produce incomplete or even misleading decisions when faced with unseen or broader operating conditions.
Federation is therefore not an optional extension but a necessary complement to FMs in these environments. Without it, long communication gaps leave local updates stale and inconsistent; corrupted or failing sensors can dominate local learning; and distribution shifts driven by seasonality, mobility, or dynamic layouts cause rapid divergence across nodes. By pooling updates from many devices, federation counterbalances these weaknesses: it reconciles delayed adaptations, dilutes the effect of corrupted inputs, and maintains alignment across diverse physical contexts. In essence, federation transforms isolated, fragile FMs into a coherent, collective intelligence that is robust to extreme environments.
This makes FFMs the best available choice for intelligent operation in harsh environments. Without the federated component, FMs risk fragmenting into narrow, unreliable local specialists; with it, they become resilient systems capable of supporting safety-critical decisions under the most demanding conditions.
III-B Challenges of FFMs in Harsh Environments
The deployment of FFMs in extreme environments is fundamentally constrained by the physical realities of the underlying infrastructure. These physical constraints translate into specific network-level challenges (depicted in Figure 5), which in turn impact the design of learning systems. In this section, we identify several network challenges that emerge from harsh environments and explain how they affect FFM training and inference.
1. Fragility of Deployed Devices and Sensors: Harsh environments impose mechanical, environmental, and chemical stress on deployed equipment. In underground mines, for instance, heavy vibrations, rockfalls, and high dust concentration frequently damage communication relays and sensors [9]. Similarly, underwater deployments face corrosion from saline water, high-pressure fluctuations, and biofouling [10], while forest environments expose devices to extreme humidity, rainfall, and wildlife interference. Frequent failures and intermittent sensing produce missing, partial, or corrupted updates. FFMs must tolerate update dropouts, filter unreliable signals, and bound the impact of noisy or adversarial contributions. Practical mechanisms include failure-aware aggregation, robust objectives (e.g., clipping or trimmed means), and confidence-weighted fusion that down-weights uncertain updates.
2. Asynchronous and Stale Updates in Harsh Networks: Many harsh environments are characterized by significant delays in data transmission due to physical medium limitations. Underwater acoustic links exhibit inherently high latency due to slow sound propagation speeds. Similarly, satellite-ground communications are constrained by orbital dynamics and long round-trip times, especially in deep space scenarios [11]. In underground and forested areas, multi-hop relays and physical obstructions introduce additional delay. The irregular contact windows and long propagation delays cause out-of-order, stale contributions that degrade global consistency. FFMs require asynchrony-tolerant protocols (e.g., buffered aggregation with age-of-update weighting), staleness-aware optimization, and contact-aware scheduling that prioritizes temporally informative updates.
Figure 5: Physical and network challenges for Federated Learning (FL) in harsh environments. Each category of harsh environment presents distinct physical constraints such as structural collapse risk, signal attenuation, and extreme humidity. These physical limitations give rise to a shared set of network-level challenges that hinder reliable FL deployment. The issues must be addressed to design robust, energy-efficient, and adaptive FL systems across diverse hostile environments.
3. Sparse and Bandwidth-Constrained Communication: In many harsh environments, both the spatial distribution of devices and the quality of communication links are severely constrained. Sparse deployments are common in deep forests, where dense vegetation and uneven terrain hinder dense sensor placement, and in underwater and satellite systems, where devices are widely spaced due to physical or operational limitations. This spatial sparsity reduces the frequency of inter-device communication and limits opportunities for collaborative learning. Compounding this challenge, the available communication links are often bandwidth-constrained and unreliable. Underwater acoustic channels offer only low-frequency, narrowband connectivity and are susceptible to multipath distortion and high bit error rates [10]. Similarly, LEO satellite links provide limited contact windows with tight bandwidth and energy budgets [12], while undermine deployments typically rely on low-rate radios. These limitations restrict both the volume and fidelity of transmitted model updates. For FFMs, such constraints necessitate the use of lightweight and highly compressed communication protocols, such as adapter-based or quantized updates, as well as the design of robust representation learning strategies that can generalize effectively from locally sparse and potentially modality-incomplete observations.
4. Intensified Data Heterogeneity and Limited Local Diversity: In harsh environments, data heterogeneity stems not only from intrinsic differences in local data distributions but also from degraded and unstable sensing conditions. While clients may operate with entirely different sensing modalities, even those using the same type of sensor can produce vastly different data due to their physical separation and exposure to varying environmental factors. For instance, two acoustic sensors deployed in different parts of an ocean may experience different levels of ambient noise or water pressure levels, resulting in significant variation in signal characteristics. Furthermore, the dynamic nature of these environments introduces rapid temporal drift, where the data distribution observed by a client can shift within hours or days due to changing weather patterns or equipment wear. Compounded by the fact that sensor nodes often have limited modality coverage and reduced sampling capacity, the sensing capability of devices in these settings is often limited due to hardware constraints, leading to reduced modality coverage and narrow local data diversity.
As a result, each client not only observes a biased slice of the global data distribution but also lacks the intra-client variety needed for stable local training. These conditions severely limit the effectiveness of traditional personalization strategies and amplify the risk of representation drift. To ensure coherent global learning, FFMs must incorporate domain-adaptive techniques and robust federated alignment mechanisms that explicitly account for both inter-client heterogeneity and sparse intra-client observations.
5. Difficulty in Node Maintenance and Replacement: Servicing deployed devices in extreme environments is often infeasible. Battery replacement for underwater or satellite nodes involves expensive and infrequent missions, while mine and forest deployments face safety and accessibility issues. As a result, nodes are expected to operate autonomously over extended periods under strict energy budgets. For FFMs, this places emphasis on designing energy-aware training schedules and communication protocols, lightweight model architectures, and mechanisms that allow models to operate reliably with limited connectivity and computational resources. Semi-decentralized learning schemes, where lower energy device-to-device communication assists the training process, can further improve sustainability in such environments [13, 14].
IV Open Research Directions
Despite the promising capabilities of FFMs in harsh and resource-constrained environments, several key challenges remain unresolved. In this section, we outline critical open problems and propose directions for future research that could further enhance the robustness, adaptability, and effectiveness of FFMs in extreme operational settings. To ground our discussion, we list in Table I several publicly available datasets spanning satellite, forest, and underwater environments that serve as benchmarks for FFMs.
Multimodal Context Alignment: In harsh and resource-constrained environments, FFMs must fuse information from multiple sensing modalities collected across heterogeneous edge devices, such as acoustic, visual, infrared, or environmental signals. In contrast to centralized foundation models, where modalities are well curated and aligned, harsh environments introduce partial modality coverage, severe fidelity variations, and missing modality co-occurrences. For example, identical sensors may degrade differently under high humidity or pressure from harsh environments, producing inconsistent fidelity across clients. Furthermore, local devices may operate with only a subset of the modalities due to limited hardware capabilities or extreme environmental limitations. This partial modality coverage can vary across clients or even over time. As a result, the embedding space can fragment, threatening semantic consistency across the network. Addressing this requires dynamic alignment mechanisms that can stitch together incomplete and noisy cross-modal links, preserve embedding coherence, and adapt to evolving modality availability under strict communication and computation constraints.
Continuous Adaptation in Evolving Conditions: The non-stationarity nature of harsh environments demands continual learning, as underlying data distributions shift due to equipment degradation, structural reconfiguration, or environmental changes. Sensor readings can fluctuate (e.g., pressure gauges drifting over time or optical sensors failing under dust) making catastrophic forgetting even more severe. FFMs must therefore adapt without erasing past knowledge, often in the absence of replay buffers (due to storage limits) or with regularizers that may fail under rapidly changing inputs. Developing robust continual learning strategies that retain knowledge while ensuring long-term adaptability remains a critical research frontier for autonomous FFMs in unpredictable and high-variability settings.
Table I: Publicly available datasets collected in harsh environments across diverse sensing modalities and tasks. These datasets span satellite, forest, and underwater settings, offering multimodal inputs such as multispectral images, point clouds, and depth maps. They support a range of downstream tasks including species identification, debris detection, and structural inspection, providing valuable benchmarks for developing and evaluating FFMs under challenging real-world conditions.
Dataset Environment Modalities Tasks
FLAIR-HUB Satellite and UAV Topographic Information, Satellite Image Time Series(SITS), VHR Land Cover Segmentation, Crop Type Segmentation
Planted Satellite SITS, VHR Forest plantation/species Identification
MADOS Satellite Multispectral Satellite Images Marine debris and oil spill mapping
FinnWoodlands Deep Forest RGB images, point clouds, depth maps Instance Segmentation, Panoptic Segmentation, Depth Completion
SUIM Underwater Image Underwater Species segmentation
LIACI Underwater Image Underwater Ship Inspection
Mitigate Performance Degradation from Drifting Effects: In FFM systems, multiple forms of “drift” can combine to undermine the accuracy and stability of the global model. In FL, data heterogeneity across clients leads to the well‐known “client drift” problem, in which each local model converges toward an optimum that fits its own data distribution, rather than the global optimum.
For FFMs, this challenge becomes even more complex due to the online learning nature of FMs. Two additional drift phenomena emerge: 1) Embedding Drift. The learned embedding gradually changes its semantic meaning over time. This can occur when the underlying data distribution evolves or when the available modality changes, making an embedding learned months ago no longer accurately represent the same concept today. 2) ID Drift. The mapping between discrete identifiers (e.g., item IDs, user IDs) and their corresponding inputs changes over time. In large‐scale systems, such as recommendation or content‐delivery platforms, old IDs frequently expire while new IDs are introduced. As a result, a given embedding vector may be reassigned to represent an entirely different entity, causing severe misalignment between model parameters and the real‐world concepts they are meant to capture. These multi-layered drifts are combined, degrading global performance and stability. Future research must design federated optimization techniques that jointly mitigate heterogeneous drifts while maintaining alignment across distributed, intermittently connected nodes.
Robustness Tuning Under Missing Communication and Corrupted Data: In harsh environments, robustness must be ensured not only against incomplete data but also against unreliable communication and sensor failures. When communication is at its weakest, or in some cases absent altogether, the global model may wait for days to receive a federated update, leaving individual clients isolated. Compounding this issue, sensors are often prone to miscalibration or degradation, producing corrupted, noisy, or even misleading signals that may be unsafe to use for model training. These conditions create severe challenges for FFMs, which must remain resilient in the face of missing updates, extreme data heterogeneity, and corrupted local inputs. Ensuring consistent and safe performance under such conditions remains an open challenge, demanding theoretical guarantees and practical mechanisms to handle both communication breakdowns and unreliable sensing sources.
Energy-Efficient and Sustainable Operation: In harsh environments where power supply is limited or intermittent, ML models must operate within strict energy constraints. Unlike conventional FM training paradigms that primarily evaluate computational cost in terms of floating-point operations per second (FLOPs), local deployment in harsh environments requires careful consideration of total energy consumption over the training and inference lifecycle. To avoid rapid battery depletion on edge devices, a joint optimization of convergence speed and energy usage becomes necessary. In this context, lightweight FM architectures, featuring compact backbones or a minimal number of trainable parameters, are especially advantageous [15]. These models reduce both local computation costs and communication overhead, which is critical given the limited bandwidth and frequent instability of communication links in such environments. Designing energy-aware training schedules and low-power inference strategies will be essential to enable real-time or periodic model updates on resource-constrained platforms, thereby supporting sustained and autonomous operation of FFMs in energy-sensitive deployment scenarios.
V Conclusion
We advocate for the development of Federated Foundation Models (FFMs) for harsh environments as a viable path forward. By combining the generalization power of foundation models with the decentralized, communication-efficient nature of federated learning, FFMs offer a framework for building intelligent systems that are both robust and adaptive in real-world adverse conditions. We have outlined the limitations of conventional methods, highlighted critical system constraints in harsh environments, and presented a vision for FFMs that balances performance, efficiency, and resilience. Looking ahead, progress depends on three directions: (i) co-design with the physical layer, where communication/scheduling protocols are designed with the need of model training; (ii) contact or event-triggered aggregation with hierarchical or semi-decentralized topologies and lightweight adapters/experts for local change without global drift; and (iii) evaluation that goes beyond accuracy to measure energy per improvement, communication per adaptation, recovery after disruption, and stability under continual drift on field datasets. FFMs in harsh environments are thus not merely bigger models distributed farther, they are engineered systems meant to remain dependable when links are unreliable and data are scarce.